# Efficient preparation of the AKLT State with Measurement-based Imaginary Time Evolution

### v2

@@ 4.2 The AKLT state preparation with qubits
Figure 5: The AKLT state preparation in the spin-$1/2$ representation for $N=3$. (a) the total state fidelity $F_{\rm tot}$ as a function of MITE subroutine round $r$. For each $r$, MITE routines are performed throughout the whole system. (b) the peak value of the amplitude function as a function of time. We perform $100$ trials of simulations, and randomly choose three of them for the illustration in this figure. The data points are recorded within the iteration for each MITE subroutine round $r$ performed throughout the whole system up to $r=100$. For all panels, we have used $N_{\rm iter}=30$, and the interaction time is chosen to be $\epsilon=0.5$. The factor $\eta$ is chosen to be $\eta=2$ such that $E_{\text{th}}=1/4$.
@@ 5 Implementation of the measurement operator
The optimization results are shown in Fig. 7, where we show both the averaged fidelity ¯$\bar{\mathcal{F}}(U,V)$ as well as the maximum fidelity $\mathcal{F}_{\rm max}(U,V)$ between the original operator $\hat{U}$ and the parametrized circuit operator $\hat{V}$ of $100$ repetitions to characterize the performance of the optimization procedure. It is found that when the circuit layer depth is larger than $4$, of all the $100$ repititions, the maximum fidelity between $\hat{U}$ and $\hat{V}$ already reaches the value larger than $99.99\%$ [Fig. 7(a)]. When the depth is larger than $6$, although $\mathcal{F}_{\rm max}(U,V)$ remains above $99.99\%$, the average fidelity ¯$\bar{\mathcal{F}}(U,V)$ starts to decrease due to overparametrization and other basin-hopping parameters chosen.
Hence in summary, we are able to approximate the unitary $\hat{U}$ with only a single-digit number of layers of $U_{3}$ and CNOT gates, where the number of CNOT gates is of the magnitude of $\sim 10$, and the fidelity between the original target weak measurement operator and the ansatz circuit is larger than $99.99\%$. Using the default transpile function from Qiskit [99] for the same measurement operator, the number of CNOT gates is 499.
Figure 6: Recompiled quantum circuit for the measurement operator (28). (left panel) The target unitary evolution operator $\hat{U}$; (right panel) the ansatz $\hat{V}$ and its quantum circuit: an initial layer composed of $U_{3}$ gates followed by an odd layer and an even layer. The odd (even) layer has $CX$ gates on odd (even) bonds followed by $U_{3}$ gates placed on the corresponding physical sites.
Figure 7: Circuit recompilation of the measurement operator (28). (a) the averaged fidelity between target operator $U$ and the ansatz circuit $V$. The average value and the error bar are calculated over $100$ repetitions. (b) the maximum of the fidelity between target operator $U$ and the ansatz circuit $V$ within every $100$ repetitions. Other parameters are: $N_{\rm iter}=100$, and $N_{\rm hop}=5$ from a basin-hopping method.
@@ 6 Discussion
We have proposed an efficient quantum algorithm based on the measurement-based imaginary time evolution (MITE) algorithm for preparation of the AKLT state. Our results show a fast convergence to the target state typically with approximately $T\sim 20$ time steps per MITE subroutine. The total number of rounds is typically $r\sim 20$. For the case of spin-$1/2$ representation, by using the variational optimization quantum circuit recompilation algorithm, the five-qubit weak measurement operator can be transformed into a circuit with $\sim 10$ CNOT gates, which allows for a direct implementation on qubit-based quantum computers. The speed of the overall convergence benefits from a optimized interaction time parameter $\epsilon$, if this is too small the convergence tends to occur much more slowly.
@@
We note that the efficiency of preparation observed here is a special case thanks to the properties of the AKLT state. There are many other interesting states that could be also prepared using the MITE approach, including exotic dissipative many-body physics [116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 18], as well as the simulation of quantum spin Hall effects beyond the solid-state systems [126, 127]. It is however an open question of whether there are other models that can be prepared with the exceptional efficiency of the AKLT state, and exactly what properties the Hamiltonian must obey to attain this efficiency. It was shown in Ref. [52] that cluster states could also be prepared efficiently due to the commuting stabilizers that define the state. Even in the generic case, the MITE approach does not require any postselection, and is therefore deterministic: one could obtain the target AKLT state for every run of the sequence. Furthermore, no knowledge of the ground state wavefunction is required for the convergence. Hence it is potentially useful as a generic procedure in a quantum simulation context, where eigenstates of Hamiltonians require preparation.
@@ Acknowledgments
T. B. is supported by the National Natural Science Foundation of China (62071301); NYU-ECNU Institute of Physics at NYU Shanghai; Shanghai Frontiers Science Center of Artificial Intelligence and Deep Learning; the Joint Physics Research Institute Challenge Grant; the Science and Technology Commission of Shanghai Municipality (19XD1423000,22ZR1444600); the NYU Shanghai Boost Fund; the China Foreign Experts Program (G2021013002L); the NYU Shanghai Major-Grants Seed Fund; Tamkeen under the NYU Abu Dhabi Research Institute grant CG008; and the SMEC Scientific Research Innovation Project (2023ZKZD55). T. C. acknowledges support from the Singapore National Research Foundation (NRF) under NRF fellowship award NRF-NRFF12-2020-0005. The computational work for this article was partially performed on resources of the National Supercomputing Centre, Singapore (NSCC) (https://www.nscc.sg), and on resources of the National University of Singapore (NUS)’s high-performance computing facilities, partially supported by NUS IT’s Research Computing group.

### v3

@@ 4.2 The AKLT state preparation with qubits
Figure 6: The AKLT state preparation in the spin-$1/2$ representation for $N=3$. (a) the total state fidelity $F_{\rm tot}$ as a function of MITE subroutine round $r$. For each $r$, MITE routines are performed throughout the whole system. (b) the peak value of the amplitude function as a function of time. We perform $100$ runs, and randomly choose three of them for the illustration in this figure, each color representing a single run. The data points are recorded within the iteration for each MITE subroutine round $r$ performed throughout the whole system up to $r=100$. For all panels, we have used $N_{\rm iter}=30$, and the interaction time is chosen to be $\epsilon=0.5$. The factor $\eta$ is chosen to be $\eta=2$ such that $E_{\text{th}}=1/4$.
@@ 5 Implementation of the measurement operator
The optimization results are shown in Fig. 8, where we show both the averaged fidelity ¯$\bar{\mathcal{F}}(U,V)$ as well as the maximum fidelity $\mathcal{F}_{\rm max}(U,V)$ between the original operator $\hat{U}$ and the parametrized circuit operator $\hat{V}$ of $100$ repetitions to characterize the performance of the optimization procedure. It is found that when the circuit layer depth is larger than $4$, of all the $100$ repetitions, the maximum fidelity between $\hat{U}$ and $\hat{V}$ already reaches the value larger than $99.99\%$ [Fig. 8(a)]. When the depth is larger than $6$, although $\mathcal{F}_{\rm max}(U,V)$ remains above $99.99\%$, the average fidelity ¯$\bar{\mathcal{F}}(U,V)$ starts to decrease due to over-parametrization and other basin-hopping parameters chosen.
Hence in summary, we are able to approximate the unitary $\hat{U}$ with only a single-digit number of layers of $U_{3}$ and CNOT gates, where the number of CNOT gates is of the magnitude of $\sim 10$, and the fidelity between the original target weak measurement operator and the ansatz circuit is larger than $99.99\%$. Using the default transpile function from Qiskit [101] for the same measurement operator, for this weak measurement operator corresponding to two spin-$1$ sites, the number of CNOT gates is 499.
Figure 7: Recompiled quantum circuit for the measurement operator in Eq. (12). (left panel) The target unitary evolution operator $\hat{U}$; (right panel) the ansatz $\hat{V}$ and its quantum circuit: an initial layer composed of $U_{3}$ gates followed by an odd layer and an even layer. The odd (even) layer has $CX$ gates on odd (even) bonds followed by $U_{3}$ gates placed on the corresponding physical sites. The operator $\hat{U}$ corresponds to two single spin-$1$ sites, where each site is transformed to two spin-$1/2$ sites as illustrated in measurement operator from Eq. (12), plus one additional ancilla qubit, thus rendering a total of $2\times 2+1=5$-qubit unitary operator.
Figure 8: Circuit recompilation of the measurement operator from Eq. (12) for $N_{M}=2$. (a) the averaged fidelity between target operator $U$ and the ansatz circuit $V$. The average value and the error bar are calculated over $100$ repetitions. (b) the maximum of the fidelity between target operator $U$ and the ansatz circuit $V$ within every $100$ repetitions. Other parameters are: $N_{\rm iter}=100$, and $N_{\rm hop}=5$ from a basin-hopping method.
@@ 6 Discussion
We have proposed an efficient quantum algorithm based on the measurement-based imaginary time evolution (MITE) algorithm for preparation of the AKLT state. Compared to MPS-based methods [78, 79], the MITE approach for preparing the AKLT state is deterministic, requiring no postselection. It addresses the limitations of MPS-based methods and reduces the preparation time to a constant with respect to the chain length $N$, despite the complexity of the quantum many-body wavefunction. Additionally, the weak measurement operator can be directly implemented on qubit-based devices. Our results show that on average, there is a fast convergence to the target state typically with approximately $T<100$ time steps per MITE subroutine. The total number of rounds is typically within $100$ for all system sizes considered. For the case of spin-$1/2$ representation, by using the variational optimization quantum circuit recompilation algorithm, the five-qubit weak measurement operator can be transformed into a circuit with $\sim 10$ CNOT gates, which allows for a direct implementation on qubit-based quantum computers. The speed of the overall convergence benefits from a optimized interaction time parameter $\epsilon$, if this is too small the convergence tends to occur much more slowly.
@@
We note that the efficiency of preparation observed here is a special case thanks to the properties of the AKLT state. There are many other interesting states that could be also prepared using the MITE approach, including exotic dissipative many-body physics [118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 18], as well as the simulation of quantum spin Hall effects beyond the solid-state systems [128, 129]. We remark that since we utilized the properties of the AKLT state for the MITE algorithm, the question of whether there are other models that can be prepared with the same remarkable efficiency as the AKLT state remains unresolved. Exploring this issue could lead to important insights into the underlying principles of quantum state preparation, particularly concerning the role of the Hamiltonian’s structure in facilitating this exceptional efficiency. Thus, a deeper investigation is needed to determine if similar models exist and what exact properties they must possess to replicate or even surpass the efficiency observed in the AKLT state. It was shown in Ref. [52] that cluster states could also be prepared efficiently due to the commuting stabilizers that define the state. Even in the generic case, the MITE approach does not require any postselection, and is therefore deterministic: one could obtain the target AKLT state for every run of the sequence. Furthermore, no knowledge of the ground state wavefunction is required for the convergence. Hence it is potentially useful as a generic procedure in a quantum simulation context, where eigenstates of Hamiltonians require preparation.
@@ Acknowledgments
T. B. is supported by the National Natural Science Foundation of China (62071301); NYU-ECNU Institute of Physics at NYU Shanghai; the Joint Physics Research Institute Challenge Grant; the Science and Technology Commission of Shanghai Municipality (19XD1423000,22ZR1444600); the NYU Shanghai Boost Fund; the China Foreign Experts Program (G2021013002L); the NYU Shanghai Major-Grants Seed Fund; Tamkeen under the NYU Abu Dhabi Research Institute grant CG008; and the SMEC Scientific Research Innovation Project (2023ZKZD55). T. C. acknowledges support from the Singapore National Research Foundation (NRF) via the Project No. NRF2021-QEP2-02-P09, and No. NRF-NRFF12-2020-0005, as well as the support by the NRF, Singapore and A*STAR under its CQT Bridging Grant. The computational work for this article was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise, and was partially performed on resources of the National Supercomputing Centre, Singapore (NSCC) (https://www.nscc.sg), and on resources of the National University of Singapore (NUS)’s high-performance computing facilities, partially supported by NUS IT’s Research Computing group.